OpenStudy (anonymous):

Last year, the JV Football team had 15 wins, 9 losses, and 1 tie. Write the ratio described in fraction form. a). Wins to Losses b). Wins to Tie c). Wins to Games

OpenStudy (anonymous):

HELP!

OpenStudy (anonymous):

a is the answer

OpenStudy (anonymous):

NO! you suppose to write the following in the fraction

OpenStudy (anonymous):

do u have option's

OpenStudy (anonymous):

i think 15 upon 10

OpenStudy (anonymous):

c= 15/25 b= 15/16 a= 15/24